OpenStudy (anonymous):

n/3 -3= 12 Please help!

OpenStudy (anonymous):

@nincompoop Can you help ??

OpenStudy (anonymous):

@mathmath333 ^^

OpenStudy (shinalcantara):

transpose first -3 to the right side of the equation n/3 = 12 + 3 n/3 = 15 cross multiply to get the value of n

OpenStudy (anonymous):

thank you SO much!!!!! @shinalcantara Can you help me with another one ?

OpenStudy (shinalcantara):

yw :) ok

OpenStudy (anonymous):

6x + 20 = -16

OpenStudy (anonymous):

@shinalcantara ^ also, would n= 45 ? for the last question ?

OpenStudy (shinalcantara):

yep

OpenStudy (shinalcantara):

for 6x + 20 - 16 just do the same concept. transpose. then this time since the it's a coefficient, divide. can you do that

OpenStudy (anonymous):

hmmm..?

OpenStudy (shinalcantara):

*6x + 20 = -16

OpenStudy (shinalcantara):

transpose 20 to the right.. or let's say subtract 20 on both sides of the equation

OpenStudy (shinalcantara):

6x + 20 - 20 = -16 -20

OpenStudy (anonymous):

-4 ?

OpenStudy (shinalcantara):

6x = -36 take note that both are negative it's like rewriting it as 6x = -16 +(-20)

OpenStudy (shinalcantara):

6x = -36 dividing both sides with 6 x = -6

OpenStudy (anonymous):

n/3-3=12 n=45

OpenStudy (anonymous):

Okay I have one more....